# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v2

@@ Appendix E Proof of Proposition 1
$\mathbb{E}[\tau]=\sum_{t=1}^{\infty}P(\tau>t)\leq\sum_{t=1}^{\infty}P\left(\log(W_{t})<\log(1/\alpha)\right),$ (42)
where $\left\{\log(W_{t})<\log(1/\alpha)\right\}$ is defined as $E_{t}$.
By the first inequality of (40), we have
$E_{t}$$\subseteq\left\{\log\left(\frac{1}{Q_{t}^{\frac{1}{2\gamma}}}\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{S_{t}^{2}}{(Q_{t}+\lvert S_{t}\rvert)}\right)\right)<\log(1/\alpha)\right\}$
$\Rightarrow E_{t}$$\subseteq\left\{{S_{t}^{2}}<\frac{4d_{*}^{2}}{2d_{*}-1}\left(Q_{t}+|S_{t}|\right)\left(\log(1/\alpha)-\log(1/Q_{t}^{\frac{1}{2\gamma}})\right)\right\}$
$\subseteq\left\{{S_{t}^{2}}<\frac{4d_{*}^{2}}{2d_{*}-1}\left(Q_{t}+\sum_{i=1}^{t}|g_{i}|\right)\left(\log(1/\alpha)-\log(1/Q_{t}^{\frac{1}{2\gamma}})\right)\right\}\quad\left(\text{Because }\left\lvert S_{t}\right\rvert=\left\lvert\sum_{i=1}^{t}g_{i}\right\rvert\leq\sum_{i=1}^{t}\left\lvert g_{i}\right\rvert\right).$ (43)
We denote $V_{t}:=\sum_{i=1}^{t}\lvert g_{i}\rvert$ and then we can get the upper bound on $V_{t}$ and $Q_{t}$ respectively. Since $\lvert g_{i}\rvert$ for any $i$ are random variables in $[0,d_{*}]$, then $V_{t}/d_{*}$ is the sum of independent random variables in $[0,1]$. By the Chernoff bound [Harvey, 2023],
$P\left(\frac{V_{t}}{d_{*}}>(1+\delta)\cdot\mathbb{E}\left[\frac{V_{t}}{d_{*}}\right]\right)\leq\exp\left(-\frac{\delta^{2}}{3}\mathbb{E}\left[\frac{V_{t}}{d_{*}}\right]\right).$ (44)
We let the right-hand side equal to $1/t^{2}$ and thus $\delta$ is $\sqrt{6\log(t)/\mathbb{E}[V_{t}/d_{*}]}$. By definition, $|g_{i}|\leq d_{i}\leq d_{*}\Rightarrow V_{t}=\sum_{i=1}^{t}|g_{i}|\leq td_{*}$. With a probability of at least $(1-1/t^{2})$, we have
$\frac{V_{t}}{d_{*}}\leq\mathbb{E}\left[\frac{V_{t}}{d_{*}}\right]+\sqrt{6\mathbb{E}\left[\frac{V_{t}}{d_{*}}\right]\cdot\log(t)}\leq t+\sqrt{6t\cdot\log(t)}\leq 2t,\quad\forall t\geq 17.$ (45)
Similarly, as for $Q_{t}=\sum_{i=1}^{t}g_{i}^{2}$, we know $g_{i}^{2}\leq d_{i}^{2}\leq d_{*}^{2}$. Then $Q_{t}/d_{*}^{2}$ is the sum of independent random variables in $[0,1]$. After applying the Chernoff bound [Harvey, 2023], we have that with a probability of at least $1-1/t^{2}$,
$\frac{Q_{t}}{d_{*}^{2}}\leq\mathbb{E}\left[\frac{Q_{t}}{d_{*}^{2}}\right]+\sqrt{6\mathbb{E}\left[\frac{Q_{t}}{d_{*}^{2}}\right]\cdot\log(t)}\leq t+\sqrt{6t\cdot\log(t)}\leq 2t,\quad\forall t\geq 17.$ (46)
Let $H_{t}=\left\{\frac{Q_{t}}{d_{*}^{2}}\leq 2t\right\}\cap\left\{\frac{V_{t}}{d_{*}}\leq 2t\right\}$. Then by (43),
$E_{t}\cap H_{t}$$\subseteq\left\{{S_{t}^{2}}<\frac{4d_{*}^{2}}{2d_{*}-1}\left(2td_{*}^{2}+2td_{*}\right)\left(\log(1/\alpha)+\log(2td_{*}^{2})^{\frac{1}{2\gamma}}\right)\right\}$
$\subseteq\left\{\left\lvert\frac{S_{t}}{d_{*}}\right\rvert<\underbrace{\sqrt{\frac{8d_{*}(d_{*}+1)t}{2d_{*}-1}\cdot\log\left((2td_{*}^{2})^{\frac{1}{2\gamma}}/\alpha\right)}}_{:=R}\right\}.$ (47)
Since $S_{t}/d_{*}=\sum_{i=1}^{t}g_{i}/d_{*}$ is the sum of independent random variables in $[-1,1]$, applying a Hoeffding bound [Harvey, 2023] gives
$P\left(\left\lvert\frac{S_{t}}{d_{*}}-\mathbb{E}\left[\frac{S_{t}}{d_{*}}\right]\right\rvert\geq u\right)\leq 2\exp\left(\frac{-u^{2}}{2t}\right).$ (48)
We still let RHS be $1/t^{2}$ to get $u=\sqrt{2t\cdot\log(2t^{2})}$. With a probability of at least $(1-1/t^{2})$ and according to the reverse triangle inequality, we have
$\left\lvert\left\lvert\frac{S_{t}}{d_{*}}\right\rvert-\left\lvert\mathbb{E}\left[\frac{S_{t}}{d_{*}}\right]\right\rvert\right\rvert\leq\left\lvert\frac{S_{t}}{d_{*}}-\mathbb{E}\left[\frac{S_{t}}{d_{*}}\right]\right\rvert\leq\sqrt{2t\cdot\log(2t^{2})}.$ (49)
This implies that,
$\left\lvert\frac{S_{t}}{d_{*}}\right\rvert\geq\left\lvert\mathbb{E}\left[\frac{S_{t}}{d_{*}}\right]\right\rvert-\sqrt{2t\cdot\log(2t^{2})}=\frac{t\Delta}{d_{*}}-\sqrt{2t\cdot\log(2t^{2})}\geq\frac{t\Delta}{d_{*}}-\sqrt{4t\cdot\log(2t)},$ (50)
where $\Delta=\lvert\mu_{x}-\mu_{y}\rvert$. The above inequality (50) is given by the fact that
$\left\lvert\mathbb{E}\left[\frac{S_{t}}{d_{*}}\right]\right\rvert$ $=\frac{\left\lvert\mathbb{E}\left[\sum_{i=1}^{t}g_{i}\right]\right\rvert}{d_{*}}$
$=\frac{\left\lvert\mathbb{E}\left[\sum_{i=1}^{t}\left(\phi({x}_{t})-\phi({y}_{t})\right)\right]\right\rvert}{d_{*}}$
$=\frac{\left\lvert\sum_{i=1}^{t}\mathbb{E}\left[\phi({x}_{t})-\phi({y}_{t})\right]\right\rvert}{d_{*}}$
$=\frac{\left\lvert t(\mu_{x}-\mu_{y})\right\rvert}{d_{*}}$
$=\frac{t\left\lvert\mu_{x}-\mu_{y}\right\rvert}{d_{*}}.$
In the following, we show $\frac{t\Delta}{d_{*}}-\sqrt{4t\cdot\log(2t)}\geq R$ for all $t\geq t_{*}$, where
$t_{*}:=\frac{32d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}\cdot\log\left(\frac{(2d_{*}^{2})^{\frac{1}{2\gamma}}\cdot 32d_{*}^{3}(d_{*}+1)\cdot t_{*}}{(2d_{*}-1)\Delta^{2}\alpha}\right),$
[…]
$\frac{t\Delta}{d_{*}}-\sqrt{4t\cdot\log(2t)}$ $\geq\sqrt{\frac{8d_{*}(d_{*}+1)t}{2d_{*}-1}\cdot\log\left(\frac{(2td_{*}^{2})^{\frac{1}{2\gamma}}}{\alpha}\right)}$
$\frac{t\Delta}{d_{*}}\geq\sqrt{4t\cdot\log(2t)}$ $+\sqrt{\frac{8d_{*}(d_{*}+1)t}{2d_{*}-1}\cdot\log\left(\frac{(2td_{*}^{2})^{\frac{1}{2\gamma}}}{\alpha}\right)}.$ (51)
Since $d_{*}>0$ ensures $\frac{8d_{*}(d_{*}+1)}{2d_{*}-1}>4$, then (F) can always hold if (52) holds.
$\frac{t\Delta}{d_{*}}$ $\geq\sqrt{\frac{16d_{*}(d_{*}+1)t}{2d_{*}-1}\cdot\log\left(\frac{(2td_{*}^{2})^{\frac{1}{2\gamma}}}{\alpha}\right)}$
$t$ $\geq{\frac{16d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}\cdot\log\left(\frac{(2td_{*}^{2})^{\frac{1}{2\gamma}}}{\alpha}\right)}.$ (52)
The derivative on both sides of the above inequality are $1$ and $\frac{16d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}\cdot\frac{1}{2\gamma t}$ separately. We want to find $t_{*}$ which can satisfy (52). If $1>\frac{16d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}\cdot\frac{1}{2\gamma t_{*}}$, then $t\geq t_{*}$ can always guarantee the original inequality (F). We first guess $t_{*}=A\cdot\log(B\cdot t_{*}^{\frac{1}{2\gamma}})$, and then let $A=\frac{32d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}$. Since $t_{*}$ is the upper bound of the RHS of (52), we have
$t_{*}$ $\geq{\frac{16d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}\cdot\log\left(\frac{(2t_{*}d_{*}^{2})^{\frac{1}{2\gamma}}}{\alpha}\right)}$
$A\cdot\log(Bt_{*}^{\frac{1}{2\gamma}})$ $\geq{\frac{16d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}\cdot\log\left(\frac{(2t_{*}d_{*}^{2})^{\frac{1}{2\gamma}}}{\alpha}\right)}$
$2\log(Bt_{*}^{\frac{1}{2\gamma}})$ $\geq\log\left(\frac{(2d_{*}^{2})^{\frac{1}{2\gamma}}}{\alpha}\cdot\frac{32d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}\right)+\log\left(\log(Bt_{*}^{\frac{1}{2\gamma}})\right).$
Since the logarithm of a logarithm grows more slowly than the logarithm itself, the term $\log\left(\log(Bt_{*}^{\frac{1}{2\gamma}})\right)$ can be neglected compared to $\log(Bt_{*}^{\frac{1}{2\gamma}})$:
$\log(Bt_{*}^{\frac{1}{2\gamma}})$ $\geq{\log\left(\frac{(2d_{*}^{2})^{\frac{1}{2\gamma}}}{\alpha}\cdot\frac{32d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}\right)}$
$\log(B)+\log(t_{*}^{\frac{1}{2\gamma}})$ $\geq{\log\left(\frac{(2d_{*}^{2})^{\frac{1}{2\gamma}}}{\alpha}\cdot\frac{32d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}\right)}.$
We denote $B=\frac{(2d_{*}^{2})^{\frac{1}{2\gamma}}\cdot 32d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}\alpha}$, the above inequality must hold for any time point $t_{*}\geq 1$. Above all, we get
$t_{*}=\frac{32d_{*}^{3}(d_{*}+1)}{(2d_{*}-1)\Delta^{2}}\cdot\log\left(\frac{(2d_{*}^{2})^{\frac{1}{2\gamma}}\cdot 32d_{*}^{3}(d_{*}+1)\cdot t_{*}}{(2d_{*}-1)\Delta^{2}\alpha}\right).$
Hence, when $t\geq t_{*}$, we have the guarantee $\frac{t\Delta}{d_{*}}-\sqrt{4t\cdot\log(2t)}\geq R$. We can further use some universal constants $C_{1}$ and $C_{2}$ to further simplify the expression. Specifically, from the above and (50), we can write
$\left\lvert\frac{S_{t}}{d_{*}}\right\rvert\geq R,\qquad\forall t\geq\frac{C_{1}\cdot d_{*}^{3}}{\Delta^{2}}\cdot\log\left(\frac{C_{2}\cdot d_{*}^{(3+\frac{1}{\gamma})}}{\Delta^{2}\alpha}\right).$ (53)
Now, by the law of total probability, for $t$ large enough such that inequalities (45), (46), and (53) all hold:
$P(E_{t})$ $=P(E_{t}\cap H_{t})+P(E_{t}\cap H_{t}^{c})$
$=P(E_{t}\cap H_{t})+P(E_{t}|H_{t}^{c})P(H_{t}^{c})$
[…]
$=\left(1-P\left(\left\lvert\frac{S_{t}}{d_{*}}\right\rvert\geq R\right)\right)+\left(1-P(H_{t})\right)$
$=\left(1-P\left(\left\lvert\frac{S_{t}}{d_{*}}\right\rvert\geq R\right)\right)+P\left(\left\{\frac{Q_{t}}{d_{*}^{2}}>2t\right\}\cup\left\{\frac{V_{t}}{d_{*}}>2t\right\}\right)\qquad\quad(\text{by definition of }H_{t})$
$\leq\left(1-P\left(\left\lvert\frac{S_{t}}{d_{*}}\right\rvert\geq R\right)\right)+P\left(\frac{Q_{t}}{d_{*}^{2}}>2t\right)+P\left(\frac{V_{t}}{d_{*}}>2t\right)$
[…]
$\leq\frac{3}{t^{2}}.$
Now we can conclude that when $t$ is large enough such that $t\geq T:=\frac{C_{1}\cdot d_{*}^{3}}{\Delta^{2}}\cdot\log\left(\frac{C_{2}\cdot d_{*}^{(3+\frac{1}{\gamma})}}{\Delta^{2}\alpha}\right)$,
$\mathbb{E}[\tau]\leq\sum_{t=1}^{\infty}P(E_{t})=T+\sum_{t\geq T}P(E_{t})\leq T+\sum_{t=T}^{\infty}\frac{3}{t^{2}}\leq T+\frac{\pi^{2}}{2}.$ (54)
The proof is now completed.
@@ Appendix F Proof of Proposition 2
Algorithm 3 Online Detecting LLMs via Online Optimization and Betting for the Composite Hypotheses Testing
@@
12:     Get a hint $d_{t+1}$ and specify the convex decision space $\mathcal{K}_{t+1}:=[-\frac{1}{2d_{t+1}},0]$.
@@
16:     Compute $\theta_{t+1}^{A}=\max\left(\min\left(\theta_{t}^{A}-\frac{1}{\gamma}\frac{z_{t}^{A}}{a_{t}^{A}},0\right),-\frac{1}{2d_{t+1}}\right)$,
17:     and compute $\theta_{t+1}^{B}=\max\left(\min\left(\theta_{t}^{B}-\frac{1}{\gamma}\frac{z_{t}^{B}}{a_{t}^{B}},0\right),-\frac{1}{2d_{t+1}}\right)$.
18:end for
19:if the source has not been declared as an LLM then
20:     Sample $Z\sim\text{Unif}(0,1)$, declare the sequence of texts $y_{t}$ is from an LLM if $W_{T}^{A}\geq 2Z/\alpha$, or $W_{T}^{B}\geq 2Z/\alpha$.
21:end if
Previously, we use the symmetry of the absolute value to get two hypothesis:
$H_{0}^{A}:\mu_{x}-\mu_{y}-\epsilon\leq 0\text{ vs. }H_{1}^{A}:\mu_{x}-\mu_{y}-\epsilon>0,$ (55)
@@
$H_{0}^{B}:\mu_{y}-\mu_{x}-\epsilon\leq 0\text{ vs. }H_{1}^{B}:\mu_{y}-\mu_{x}-\epsilon>0.$ (56)
We now choose a nonpositive $\theta_{t}\in[-1/2d_{t},0]$ to ensure the property of nonnegative supermartingale wealth under $H_{0}$ and a fast wealth growth under $H_{1}$, rather than maintaining the same range as the original hypothesis in Chugg et al. [2023]. The wealth now becomes
$W_{t}^{A}=W_{0}\cdot\prod_{i=1}^{t}\left(1-\theta_{t}(g_{t}-\epsilon)\right)=\prod_{i=1}^{t}\left(1-\theta_{t}(g_{t}-\epsilon)\right)$ (57)
@@
$W_{t}^{B}=W_{0}\cdot\prod_{i=1}^{t}\left(1-\theta_{t}(-g_{t}-\epsilon)\right)=\prod_{i=1}^{t}\left(1-\theta_{t}(-g_{t}-\epsilon)\right)$ (58)
The parameter $\epsilon$ is a small positive constant, and the upper bound of the text score difference ($d_{t}\geq\lvert\phi(x_{t})-\phi(y_{t})\rvert$), is always set conservatively large, which ensures that $d_{t}\geq\epsilon$. Thus, the range of $\theta_{t}(g_{t}-\epsilon)$ and $\theta_{t}(-g_{t}-\epsilon)$, which is $[-(d_{t}-\epsilon)/2d_{t},(d_{t}+\epsilon)/2d_{t}]$, will always fall within the interval $[-1,1]$. It continues to satisfy the requirement for the wealth to remain nonnegative since the constant $\epsilon>0$ is always small. We can decouple $W_{t}$ into $W_{t}^{A}$ and $W_{t}^{B}$ to preserve their properties of supermartingales. Take $W_{t}^{A}$ as an example, under the corresponding null hypothesis $H_{0}^{A}:\mu_{x}-\mu_{y}\leq\epsilon$, i.e., $\mathbb{E}_{P}\left[\phi({x}_{t})-\phi({y}_{t})\right]\leq\epsilon$ for $P\in H_{0}^{A}$. We now select non-positive fractions $\theta_{t}\leq 0$ and the payoff $S_{t}^{A}=1-\theta_{t}(\phi({x}_{t})-\phi({y}_{t})-\epsilon)$, then
$\mathbb{E}_{P}[W_{t}^{A}|\mathcal{F}_{t-1}]=\mathbb{E}_{P}\left[W_{t-1}^{A}\times S_{t}^{A}\Bigg{|}\mathcal{F}_{t-1}\right]=W_{t-1}^{A}\left(1-\theta_{t}\cdot\left(\mathbb{E}_{P}[\phi({x}_{t})-\phi({y}_{t})]-\epsilon\right)\right)\leq W_{t-1}^{A}.$ (59)
As for the other null hypothesis $H_{0}^{B}:\mu_{y}-\mu_{x}\leq\epsilon$, i.e., $\mathbb{E}_{P}\left[\phi({y}_{t})-\phi({x}_{t})\right]\leq\epsilon$, we can get the same result. The Ville’s inequality again gives
$P(\exists t\leq T:W_{t}^{A}\geq 2/\alpha)\leq\alpha/2,$ (60)
@@
$P(\exists t\leq T:W_{t}^{B}\geq 2/\alpha)\leq\alpha/2.$ (61)
@@
$P(\exists t\leq T:(W_{t}^{A}\geq 2/\alpha)\cup(W_{t}^{B}\geq 2/\alpha))\leq\alpha$ (62)

### v3

@@ Appendix E Proof of Proposition 1
$=\mathbb{P}\left[\ln(W_{t}(\theta_{*}))-\text{Regret}_{t}(\theta_{*})-\mathbb{E}[\ln(W_{t}(\theta_{*}))]<\ln\left(\frac{1}{\alpha}\right)-\mathbb{E}[\ln(W_{t}(\theta_{*}))]\right]$
$=\mathbb{P}\left[\ln(W_{t}(\theta_{*}))-\mathbb{E}[\ln(W_{t}(\theta_{*}))]<\text{Regret}_{t}(\theta_{*})+\ln\left(\frac{1}{\alpha}\right)-t\omega_{*}\right].$ (21)
Now we are going to show that $\text{Regret}_{t}(\theta_{*})+\ln\left(\frac{1}{\alpha}\right)-t\omega_{*}\leq\frac{-t\omega_{*}}{2}$ when $t$ is sufficiently large. Then, from equation 21, we will have $\mathbb{P}\left[\ln(W_{t}(\theta_{*}))-\mathbb{E}[\ln(W_{t}(\theta_{*}))]<\text{Regret}_{t}(\theta_{*})+\ln\left(\frac{1}{\alpha}\right)-t\omega_{*}\right]\leq\mathbb{P}\left[\ln(W_{t}(\theta_{*}))-\mathbb{E}[\ln(W_{t}(\theta_{*}))]<-\frac{t}{2}\omega_{*}\right]$.
From the regret bound of ONS equation 15, it suffices to show that
$\frac{1}{2\gamma}\ln\left(1+\sum_{s=1}^{t}\frac{g_{s}^{2}}{(1-g_{s}\theta_{s})^{2}}\right)+\frac{\gamma}{2}D_{1}^{2}+\ln\left(\frac{1}{\alpha}\right)\leq\frac{t}{2}\omega_{*}.$ (22)
We will require the time $t$ to satisfy
$t\geq\frac{2\gamma D_{1}^{2}}{\omega_{*}},$ (23)
which leads to $\frac{\gamma}{2}D_{1}^{2}\leq\frac{t\omega_{*}}{4}$. Therefore, to guarantee equation 22, it suffices to find $t$ such that
$\frac{t\omega_{*}}{4}\geq\ln\left(\frac{1}{\alpha}\right)+\frac{1}{2\gamma}\ln\left(1+\sum_{s=1}^{t}\frac{g_{s}^{2}}{(1-g_{s}\theta_{s})^{2}}\right).$
We further note that $\frac{g_{s}^{2}}{(1-g_{s}\theta_{s})^{2}}\leq\frac{(d_{*})^{2}}{(1-1/2)^{2}}=4(d_{*})^{2}$, which in turns implies a sufficient condition:
$\frac{t\omega_{*}}{4}\geq\ln\left(\frac{1}{\alpha}\right)+\frac{1}{2\gamma}\ln\left(1+4t(d_{*})^{2}\right).$
The above condition together with equation 23 suggests that $\text{Regret}_{t}(\theta_{*})+\ln\left(\frac{1}{\alpha}\right)-t\omega_{*}\leq\frac{-t\omega_{*}}{2}$ holds when
$t\gtrsim\frac{1}{\omega_{*}}\ln\left(\frac{(d_{*})^{1/\gamma}}{\omega_{*}\alpha}\right).$ (24)
Under equation 24, we hence have
$\mathbb{P}\left[W_{t}<\frac{1}{\alpha}\right]$ $\leq\mathbb{P}\left[\ln(W_{t}(\theta_{*}))-\mathbb{E}[\ln(W_{t}(\theta_{*}))]<-\frac{t}{2}\omega_{*}\right].$ (25)
Denote $\psi_{t}:=\ln(1-g_{t}\theta_{*})-\mathbb{E}[\ln(1-g_{t}\theta_{*})]\in[\ln(1/2)-\omega_{*},\ln(3/2)-\omega_{*}]$. We have that $\ln(W_{t}(\theta_{*}))-\mathbb{E}[\ln(W_{t}(\theta_{*}))]=\sum_{s=1}^{t}\psi_{s}$, which is a sum of zero-mean i.i.d bounded random variables. Specifically, $\psi_{t}$ is a sub-Gaussian with parameter $\sigma:=\frac{1}{2}\ln(3)$. By Hoeffding’s inequality, we have
$\mathbb{P}\left[\frac{1}{t}\sum_{s=1}^{t}\psi_{s}\leq-c\right]\leq\exp\left(-\frac{tc^{2}}{2\sigma^{2}}\right),$ (26)
for any constant $c>0$. Then,
$\mathbb{P}\left[\ln(W_{t}(\theta_{*}))-\mathbb{E}[\ln(W_{t}(\theta_{*}))]<-\frac{t}{2}\omega_{*}\right]=\mathbb{P}\left[\frac{1}{t}\sum_{s=1}^{t}\psi_{s}<-\frac{1}{2}\omega_{*}\right]\leq\exp\left(-\frac{(\omega_{*})^{2}}{2(\ln 3)^{2}}t\right),$
where we let $c\leftarrow\frac{\omega_{*}}{2}$ in equation 26. Hence, we obtain
$\mathbb{P}\left[W_{t}<\frac{1}{\alpha}\right]\leq\exp\left(-\frac{(\omega_{*})^{2}}{2(\ln 3)^{2}}t\right).$ (27)
Let $H_{t}$ be the event that we stop at time $t$. Then,
$\mathbb{P}[t=\infty]$ $=\mathbb{P}\left[\lim_{t\to\infty}\cap_{s\leq t}\neg H_{s}\right]$
$=\lim_{t\to\infty}\mathbb{P}\left[\cap_{s\leq t}\neg H_{s}\right]$
$\leq\lim_{t\to\infty}\mathbb{P}\left[\neg H_{t}\right]$
$=\lim_{t\to\infty}\mathbb{P}\left[W_{t}<\frac{1}{\alpha}\right]$
$\leq\lim_{t\to\infty}\exp\left(-\frac{(\omega_{*})^{2}}{2(\ln 3)^{2}}t\right)$
$=0,$
where the last equality is by equation 27. This shows that the test power is one.
3. Expected Stopping Time.
Now we upper-bound the expected stopping time as follows. Denote $t_{*}$ the time when equation 25 starts to hold at any $t\geq t_{*}$. We have
$\mathbb{E}[\tau]$ $=\sum_{t=1}^{\infty}\mathbb{P}\left[\tau>t\right]$
$=\sum_{t=1}^{\infty}\mathbb{P}\left[\cap_{s\leq t}\neg H_{s}\right]$
$\leq\sum_{t=1}^{\infty}\mathbb{P}\left[\neg H_{t}\right]$
$=\sum_{t=1}^{\infty}\mathbb{P}\left[W_{t}<\frac{1}{\alpha}\right]$
$\leq t_{*}+\sum_{t=1}^{\infty}\exp\left(-\frac{(\omega_{*})^{2}}{2(\ln 3)^{2}}t\right)$
$\leq t_{*}+\frac{1}{\exp\left(\frac{(\omega_{*})^{2}}{2(\ln 3)^{2}}t\right)-1}$
$\leq t_{*}+\frac{2(\ln 3)^{2}}{\omega_{*}^{2}},$ (28)
where the last inequality uses $\exp(z)\geq 1+z$ for any $z\geq 0$.
To proceed, we recall the notation $\omega_{*}:=\mathbb{E}\left[\ln(1-g\theta_{*})\right]$. We now let the benchmark $\theta_{*}$ to be the one that maximizes the expected wealth in one round, i.e., $\mathbb{E}\left[\ln(1-g\theta)\right]$, in the decision space $\mathcal{K}_{*}$.
By the inequality $\ln(1+c)\geq c-c^{2}$ for any $|c|\leq\frac{1}{2}$, we have
$\omega_{*}=\max_{\theta\in K_{*}}\mathbb{E}\left[\ln\left(1-g\theta\right)\right]$ $\geq\max_{\theta\in K_{*}}\mathbb{E}\left[-\theta g-\theta^{2}g^{2}\right]$
$=\max_{\theta\in K_{*}}-\theta(\mu_{x}-\mu_{y})-\theta^{2}\left(\mathrm{Var}[g]+(\mu_{x}-\mu_{y})^{2}\right),$
$\approx\frac{\Delta^{2}}{\mathrm{Var}[g]+\Delta^{2}}$
$\approx\frac{\Delta^{2}}{d_{*}^{2}+\Delta^{2}},$
where we note $\mathbb{E}[g]=\mathbb{E}[\phi(x)-\phi(y)]=\mu_{x}-\mu_{y}$, $\mathrm{Var}[g]=\mathbb{E}[g^{2}]-(\mathbb{E}[g])^{2}$, and $\mathrm{Var}[g]\leq d_{*}^{2}$.
Combining equation 24 and equation 28, together with the expression of $\omega_{*}$ above, we have
$\mathbb{E}[\tau]=\mathcal{O}\left(\frac{d_{*}^{2}}{\Delta^{2}}\ln\left(\frac{d_{*}}{\alpha\Delta}\right)+\frac{d_{*}^{4}}{\Delta^{4}}\right).$ (29)
@@ Appendix F Proof of Proposition 2
Algorithm 3 Online Detection of LLMs via Online Optimization and Betting for the Composite Hypotheses Testing
@@
12:     Get a hint $d_{t+1}$ and specify the decision space $\mathcal{K}_{t+1}:=[-\frac{1}{2d_{t+1}},0]$.
@@
16:     Compute $\theta_{t+1}^{A}=\max\left(\min\left(\theta_{t}^{A}-\frac{1}{\gamma}\frac{z_{t}^{A}}{a_{t}^{A}},0\right),-\frac{1}{2d_{t+1}}\right)$, $\theta_{t+1}^{B}=\max\left(\min\left(\theta_{t}^{B}-\frac{1}{\gamma}\frac{z_{t}^{B}}{a_{t}^{B}},0\right),-\frac{1}{2d_{t+1}}\right)$.
17:end for
18:if the source has not been declared as an LLM then
19:     Sample $Z\sim\text{Unif}(0,1)$, declare the sequence of texts $y_{t}$ is from an LLM if $W_{T}^{A}\geq 2Z/\alpha$ or $W_{T}^{B}\geq 2Z/\alpha$.
20:end if
Recall we formulated two one-sided hypotheses:
$H_{0}^{A}:\mu_{x}-\mu_{y}-\epsilon\leq 0\text{ vs. }H_{1}^{A}:\mu_{x}-\mu_{y}-\epsilon>0,$
@@
$H_{0}^{B}:\mu_{y}-\mu_{x}-\epsilon\leq 0\text{ vs. }H_{1}^{B}:\mu_{y}-\mu_{x}-\epsilon>0.$
1. Level-$\alpha$ Sequential Test. In contrast to the betting interval employed in Chugg et al. [2023] for composite hypotheses test, which includes both positive and negative values, we specify a nonpositive decision space $\theta_{t}\in[-1/2d_{t},0]$ in order to ensure that the wealth process is a supermartingale under $H_{0}$, while still enabling fast wealth growth under the alternative $H_{1}$. The resulting wealth process is now given by:
$W_{t}^{A}=W_{0}\cdot\prod_{s=1}^{t}\left(1-\theta_{s}(g_{s}-\epsilon)\right)=\prod_{s=1}^{t}\left(1-\theta_{s}(g_{s}-\epsilon)\right)$
@@
$W_{t}^{B}=W_{0}\cdot\prod_{s=1}^{t}\left(1-\theta_{s}(-g_{s}-\epsilon)\right)=\prod_{s=1}^{t}\left(1-\theta_{s}(-g_{s}-\epsilon)\right),$
where $g_{s}=\phi(x_{s})-\phi(y_{s})$. Recall that $\epsilon$ is a nonnegative constant satisfying $\epsilon\leq d_{t}$ for all $t$. Consequently, the range of $\theta_{t}(g_{t}-\epsilon)$ and $\theta_{t}(-g_{t}-\epsilon)$, given by $[-(d_{t}-\epsilon)/2d_{t},(d_{t}+\epsilon)/2d_{t}]$, always lies within the interval $[-1,1]$. This guarantees that the wealth process remains nonnegative. Moreover, the supermartingale property of the wealth processes $W_{t}^{A}$ and $W_{t}^{B}$ can still be preserved. Take $W_{t}^{A}$ as an example, under the corresponding null hypothesis $H_{0}^{A}:\mu_{x}-\mu_{y}\leq\epsilon$, i.e., $\mathbb{E}_{P}\left[\phi({x}_{t})-\phi({y}_{t})\right]\leq\epsilon$ for $P\in H_{0}^{A}$. Then,
$\mathbb{E}_{P}[W_{t}^{A}|\mathcal{F}_{t-1}]=\mathbb{E}_{P}\left[W_{t-1}^{A}\left(1-\underbrace{\theta_{t}}_{\leq 0}\cdot\left(\phi({x}_{t})-\phi({y}_{t})-\epsilon\right)\right)\Bigg{|}\mathcal{F}_{t-1}\right]\leq W_{t-1}^{A},$
where the inequality is because $\theta_{t}$ is fully determined by $\mathcal{F}_{t-1}$, and hence, given $\mathcal{F}_{t-1}$, $\theta_{t}$ is independent of $\left(\phi({x}_{t})-\phi({y}_{t})-\epsilon\right)$, where the latter has conditional expectation less than or equal to zero under $H_{0}^{A}:\mu_{x}-\mu_{y}\leq\epsilon$.
As for the other null hypothesis $H_{0}^{B}:\mu_{y}-\mu_{x}\leq\epsilon$, i.e., $\mathbb{E}_{P}\left[\phi({y}_{t})-\phi({x}_{t})\right]\leq\epsilon$, it can be shown that the wealth process $W_{t}^{B}$ also forms a nonnegative supermartingale.
Applying the randomized Ville’s inequality gives us
$P(\exists t\geq 1:W_{t}^{A}\geq 2/\alpha)\leq\alpha/2,$ (30)
@@
$P(\exists t\geq 1:W_{t}^{B}\geq 2/\alpha)\leq\alpha/2.$ (31)
@@
$P(\exists t\geq 1:(W_{t}^{A}\geq 2/\alpha)\cup(W_{t}^{B}\geq 2/\alpha))\leq\alpha,$